# iSeg: An Iterative Refinement-based Framework for Training-free Segmentation

### v2

@@ 5 Conclusion
In this paper, we propose an iterative refinement framework for training-free segmentation using pre-trained stable diffusion model, named iSeg. Our proposed iSeg introduces two novel modules: entropy-reduced self-attention module and category-enhanced cross-attention module. The entropy-reduced self-attention module aims to suppress irrelevant global information within self-attention map, while the category-enhanced cross-attention module pays more attention to the features of given segmented categories. With these two proposed modules, our proposed method can generate accurate segmentation masks without any training on segmentation datasets. We perform the experiments on various datasets and segmentation tasks to demonstrate the efficacy of our proposed method.
@@
Appendix A Algorithms
Algorithm 1 presents the pseudo code about how to refine the cross-attention by entropy-reduced self-attention iteratively using our proposed iSeg.
Appendix B More implementation details
Here we present more implementation details of our proposed method on different segmentation tasks, including weakly-supervised semantic segmentation, open-vocabulary semantic segmentation, unsupervised segmentation, and mask generation for synthetic dataset.
Weakly-supervised semantic segmentation. In weakly-supervised semantic segmentation, one of the key is to generate pseudo pixel-level masks for training set. We compare our training-free method with existing approaches for pixel-level mask generation on two datasets, including PASCAL VOC 2012 [68] and MS COCO 2014 [69]. Specifically, we use the categories existing in image to generate a template such as ‘A photograph of [CLS A] and [CLS B] and other objects and background’. Here we compose the sentence with some general words, such as ‘object’ and ‘background’ rather than specific background category associated with the specific dataset similar to CLIPES [7].
Open-vocabulary semantic segmentation. Open-vocabulary semantic segmentation aims to perform segmentation of arbitrary categories. We compare our training-free method with existing approaches on three datasets, including PASCAL VOC 2012 (VOC), PASCAL-Context (Context), and MS COCO Object (Object). To perform open-vocabulary semantic segmentation, we employ the pre-trained CLIP to generate image-level labels as text prompts following TagCLIP [67]. Afterwards, we keep the same settings as weakly-supervised semantic segmentation task to obtain the final mask.
Unsupervised segmentation. We compare our training-free method with other unsupervised segmentation approaches on two datasets, including COCO-Stuff-27 [74] and Cityscapes [73]. We integrate our proposed method into the training-free approach DiffSeg [15] as a post-processing. Specifically, we treat the segmentation results of DiffSeg as initial cross-attention maps, and employ our proposed entropy-reduced self-attention module to perform iterative refinement, while the other settings are consistent with DiffSeg.
Mask generation for synthetic dataset. Recently, some researchers have explored generating synthetic dataset for semantic segmentation. In this task, it is important to generate the mask labels for synthetic images. DiffuMask [18] utilizes cross-attention maps in diffusion model to generate pseudo masks for synthetic images, which are further refined by AffinityNet. Here, we can directly apply our iterative refinement to improve the initial pseudo masks generated by DiffuMask on synthetic dataset without requiring any additional refinements like AffinityNet.
Appendix C Qualitative results
Cross-attention map after different iterations. Fig. 8 visualizes some refined cross-attention maps of our proposed iSeg at different refinement iterations. With the increment of iterations ranging from 0 to 8, our proposed iSeg can generate more accurate cross-attention maps of corresponding objects.
Results on synthetic images. Fig. 9 presents some segmentation results of our proposed iSeg on synthetic images generated by DiffuMask [18]. It can be observed that our proposed method can generate accurate segmentation masks for diverse synthetic images.
Results on natural images. Fig. 10 presents some segmentation results of our proposed iSeg on PASCAL VOC 2012 and MS COCO. It can be observed that our proposed method can generate accurate segmentation masks for various categories.
Results of unsupervised segmentation. Fig. 11 presents some unsupervised segmentation results of our proposed iSeg on COCO-Stuff-27 datasets. Our proposed method is able to accurately group the pixels into different objects.
Cross-domain results. Fig. 12 presents some kind of image segmentation on DomainNet [75]. Our proposed method is able to accurately segment different kinds of objects including sketch, painting and clipart according to given category.
Results using different interactions. Fig. 13 presents more segmentation results using different interactions, such as points, lines, and boxes.

### v3

@@ 5 Conclusion
In this paper, we propose an iterative refinement framework for training-free segmentation using pre-trained stable diffusion model, named iSeg. Our iSeg introduces two novel modules: entropy-reduced self-attention module and category-enhanced cross-attention module. The entropy-reduced self-attention module aims to suppress irrelevant global information within self-attention map, while the category-enhanced cross-attention module pays more attention to the features of given segmented categories. With these two proposed modules, our proposed method can generate accurate segmentation masks without any training on segmentation datasets. We perform the experiments on various datasets and segmentation tasks to demonstrate the efficacy of our proposed method.